### v1

# Conventional Quantum Theory Does Not Support A Coherent Relational Account

ABSTRACT. I review a counterexample to the frequent claim that discrepancies among observers resulting from conventional quantum theory’s inability to define “measurement,” such as those arising in the Wigner’s Friend thought experiment, remain private and incommensurable. I consider the implications for a recent attempt to shield Relational Quantum Mechanics from such inconsistencies and conclude that it is not successful.
1. Introduction and Background
This paper reviews the challenge to conventional quantum theory (which I will henceforth call CQM) from inconsistencies arising from its prediction of macroscopic superpositions such as Schrödinger’s Cat (SC) and Wigner’s Friend (WF). I focus on an interpretational approach first proposed by Rovelli (1996) called “Relational Quantum Mechanics”(henceforth RQM). RQM is revisited with modifications in Adlam and Rovelli 2022 (henceforth AR2022), The latter paper is the subject of the present critique.
The topic at hand is therefore fundamentally the measurement problem (MP) of CQM, since thought experiments involving ‘absurd’ macroscopic superpositions (the characterization in Frauchiger and Renner 2018, and clearly Schrödinger’s intent) are all illustrations of the MP. It is important to keep that point in mind in what follows. For clarity, CQM is the standard approach to quantum theory that denies any real physical non-unitarity and therefore must treat observed determinate outcomes as brute empirical facts with no physical explanation within the theory (more strongly, observed outcomes fail to be predicted by the theory, as the MP macroscopic superposition paradoxes illustrate). This point is granted by AR2022, who say (in the context of a WF experiment):
“Unitary quantum mechanics does not provide any mechanism for a single measurement outcome to be selected and actualised for Alice in the first place, so it certainly cannot tell us anything about the relationship between her outcome and Bob’s outcome - this is a question which lies entirely outside the unitary part of the theory…” (AR2022)
The locution ‘unitary part of the theory’ might seem to suggest that there is another ‘part of the theory,’ but that merely leaves room for an ad hoc application of a ‘collapse postulate’ notoriously subject to the ‘Shifty Split’ between the system to which CQM is taken as applying and some putative ‘outside observer’. This situation is what gives rise to the Wigner’s Friend and related inconsistencies such as that of the Frauchiger-Renner Paradox (Frauchiger and Renner, 2018). Of course, there are many attempts to deal with the MP and resulting inconsistencies arising from CQM and its Shifty Split. The paper critiqued herein, AR2022, is one of those attempts; I argue in what follows that it fails.
In contrast to CQM, formulations involving objective reduction or collapse (henceforth OR) [^1] Some examples are the gravitational collapse model of Penrose 2000 and the Transactional Interpretation (Cramer 1986, most recent relativistic version in Kastner 2022) are not subject to the MP and thus do not lead to macroscopic superpositions such as that of Bob and his lab in the Wigner’s Friend paradox. This is the case regardless of one’s opinion of the theoretical or practical adequacy of such theories, which is a separate issue. Interestingly, one of the authors of AR2022 (EA) does acknowledge that a particular OR formulation solves the MP in a footnote to a paper entitled “Do We Have Any Viable Solution to the Measurement Problem?”, where she says:
“…the transactional interpretation…is a solution to the measurement problem which has the property of being retrocausal.” (Adlam 2023). [^2] Adlam’s references are out of date, however; in addition to Cramer (1986), a more recent, fully relativistic account is Kastner (2022); see also Kastner (2018) in regards to TI’s solution to the MP.
(This comment was made as an aside in the context of properties a particular formulation or interpretation might have. Oddly, however, although acknowledged by the author as a solution to the MP in the quoted footnote, the transactional interpretation is not mentioned anywhere in the body of the paper, despite its title—even with the emphasis on “Any”). [^3] This phenomenon, in which an author disregards a formulation that they themselves admit in a footnote solves the problem that is the subject of their paper, would provide for an interesting study in the sociology and psychology of science. However, such considerations are beyond the scope of this paper.
Thus, at the risk of repetition (for some emphasis seems necessary), OR formulations eliminate the MP so that such ‘paradoxes’ as Schrödinger’s Cat and Wigner’s Friend do not arise in the first place. It is important to keep this point—that the notorious observer-related paradoxes do not arise under OR formulations—firmly in mind, since it appears to have been somehow forgotten in AR2022. To elaborate a bit, although it should be obvious: OR formulations involve genuine physical non-unitarity in some form, such that the measurement process is defined at the appropriate micro-level and an interaction constituting a ‘measurement’ according to that theory results in an objectively existing outcome such that the measured system is accurately described by the corresponding eigenstate. This means that an OR formulation (such as the transactional interpretation, TI and its relativistic version RTI; cf. Kastner 2022) would dictate that Schrödinger’s Cat is not in a macroscopic superposition, since there definitely is or is not an atomic decay in the box at some particular time (cf. Kastner and Cramer 2018). It would similarly dictate that Bob and his lab are not in a macroscopic superposition under the envisioned circumstances of Bob’s measurement. Thus Wigner, under an OR theory, will accordingly predict that his friend Bob will actually get an outcome corresponding to an eigenstate for his measured system, and Wigner will assign the appropriate proper mixed state to Bob’s system if Wigner does not currently know that outcome. Wigner will not describe Bob and his lab by a macroscopic superposition, for the OR theory would not predict that (as discussed above). Only CQM predicts that. As noted above, whether or not one happens to like or accept or believe a particular OR theory to be fully adequate is irrelevant here: such a theory specifically says that measurements involve collapse and predicts that superpositions therefore do not persist beyond that point. And indeed Adlam acknowledges in (2023) that one such OR theory (the Transactional Interpretation)“solves the measurement problem,” as noted above. So one cannot purport to be applying an OR theory while assuming the existence of macroscopic superpositions denied by OR theories. Yet, as we will see in what follows, AR2022 do just that.
2. Critique
The above quote from AR2022, namely: “Unitary quantum mechanics does not provide any mechanism for a single measurement outcome to be selected and actualised for Alice in the first place….” illustrates the conventional assumption of CQM that the actual physical content of quantum theory must be restricted to unitary evolution, and therefore that measurement outcomes are not accounted for by the theory but are brute matters of fact for which some ‘interpretational’ account is needed. However, before one even gets to issues of consistency or correspondence, it is worth noting that RQM simply helps itself to the observed fact, admitted by the authors as not at all accounted for by CQM, that measurements have outcomes—even if only ‘relative’ to some observer or another. This is not a matter of the theory being ‘interpreted’ but merely an ad hoc patching of its key, acknowledged lacuna, since the theory provides for no outcomes, ‘relative’ or otherwise. In this sense, RQM is not really an interpretation of CQM but a different theory consisting of CQM along with an ad hoc postulate in place of the projection postulate; namely “outcomes occur relative to observers.” The latter is not a physical postulate that serves to explain empirical data, but rather simply postulates the empirical data that its underlying theory admittedly fails to explain. RQM thus postulates the explanandum, and that makes RQM vacuous as an explanans.
In contrast, OR theories, as noted above, provide for the actualization of an outcome under specific physical circumstances involving some form of non-unitarity, and are not observer-dependent; thus in OR theories there is never an ‘inner’ or ‘outer’ observer’ as in the WF scenario. There is a well-defined point at which collapse/reduction occurs as an invariant physical event, and any observer describes the situation accordingly. However, CQM is subject to the WF and related experiments, in which an ‘inner observer’ and ‘outer observer’ disagree on the states of the systems involved, which raises the possibility, made explicit in the Frauchiger-Renner paradox, that the ‘inner observer’ could be found to make a ‘wrong’ prediction. The latter circumstance refutes the central postulate of RQM, stated in Laudisa and Rovelli (2021) as follows:
“The physical assumption at the basis of RQM is the following postulate: The probability distribution for (future) values of variables relative to $S$ depend on (past) values of variables relative to $S$ but not on (past) values of variables relative to another system $S^{\prime}$.” [^4] Laudisa and Rovelli (2021). It should be noted, however, that this is not a “physical assumption.” It is a calculational, instrumentalist postulate that says nothing about any underlying physics. In contrast, Boltzmann’s assumption that the large-scale laws of thermodynamics were based on the microscopic behaviors of atoms was a physical postulate.
The form of RQM reviewed and slightly revised in AR2022—call it RQM22—thus abandons the central postulate of RQM (although the present author has been unable to find any admission on the part of RQM advocates of the fact that its core assumption is violated by the CQM that they purport to be interpreting). The original theoretically contradicted postulate is now replaced by the following five postulates:
1. Relative facts: Events, or facts, can happen relative to any physical system. No hidden variables: unitary quantum mechanics is complete
2. Relations are intrinsic: the relation between any two systems A and B is independent of anything that happens outside these systems’ perspectives
3. Relativity of comparisons: it is meaningless to compare the accounts relative to any two systems except by invoking a third system relative to which the comparison is made.
4. Measurement: an interaction between two systems results in a correlation within the interactions between these two systems and a third one; i.e. with respect to a third system W, the interaction between the two systems S and F is described by a unitary evolution that potentially entangles the quantum states of S and F.
5. Internally consistent descriptions: In a scenario where F measures S, and W also measures S in the same basis, and W then interacts with F to ‘check the reading’ of a pointer variable (i.e. by measuring F in the appropriate ‘pointer basis’), the two values found are in agreement. (quoted from AR2022).
These postulates introduce undefined terms such as the ‘fact’ and ‘event’ of Postulate 1. As mentioned above, if these are supposed to denote measurement outcomes, the authors help themselves to the occurrence of outcomes ‘for observers’ despite the fact, acknowledged by them, that “Unitary quantum mechanics does not provide any mechanism for a single measurement outcome to be selected and actualised for Alice in the first place…”. While AR2022 go on to talk about “flashes” as a matter of purported ontology, under CQM these ad hoc entities can never constitute the “observer-independent events” offered, as we will see in what follows. In addition, Postulate 3’s purpose seems unclear. If the intent is to rule out an observer-independent account, one might wonder why that would be necessary in view of the authors’ claim that their interpretation yields observer-independent facts. [^5] Deeming certain questions “meaningless” can serve to evade legitimate questions; in particular, comparison of accounts would seem to be meaningful if one is purporting to provide a physical and even ‘observer-independent’ ontology of the world, which AR2022 are doing.
As an aside, the present author certainly agrees with the authors that one should aim for a fully relativistic interpretation, one which takes into account that spacetime could be emergent from the quantum level and thus is not assumed as a fixed background or container for events, as in their remark: “there is an alternative view which suggests that in fact, spacetime should be understood to emerge from a background of quantum events.” Indeed a comprehensive account of just how this could work is provided in Schlatter and Kastner (2023). The quote from Born is also apt: “One should not transfer the concept of space-time as a four-dimensional continuum from the macroscopic world of common experience to the atomistic world; manifestly the latter requires a different type of manifold.” However, these considerations do not save RQM, nor its recent reworking in AR2022, from the problems discussed herein.
The authors review the Wigner’s Friend scenario to which CQM is subject, in which an ‘inner observer’ (Bob) performs a measurement M, finds an outcome and attributes the relevant eigenstate to his system, while Wigner (the ‘outer observer’) describes Bob with a coherent superposition, so that the two disagree on what quantum theory predicts. They claim that this disagreement cannot be corroborated in any way: “Wigner and Bob simply live within incommensurate realities now, and no attempt to reach across and bridge the gap can possibly succeed.”. (AR2022)
The authors then focus on this alleged situation (which is falsified by a counterexample to be discussed in Section 2b) as something to be improved upon and to that end suggest replacing Postulate 4 with Postulate 4.1 involving “cross-perspective links”:
“Definition 4.1. Cross-perspective links: In a scenario where some observer Alice measures a variable V of a system S, then provided that Alice does not undergo any interactions which destroy the information about V stored in Alice’s physical variables, if Bob subsequently measures the physical variable representing Alice’s information about the variable V, then Bob’s measurement result will match Alice’s measurement result.” (AR20220)
However, 4.1 does not succeed in remedying the essential inconsistency faced by CQM, as we will see in what follows. We now elucidate four consequential errors made by the authors:
Error 1: Failure to actually apply an OR theory while claiming to apply it
Error 2: Failure to acknowledge empirically consequential nature of inconsistent predictions
Error 3: Unsupported claim that Wigner’s measurement ‘destroys’ Bob’s outcome
Error 4: Infinite regress involving ‘superobservers’ nullifies consistency claim
I now discuss each of these errors in turn.
2a. Error 1: Failure to apply an OR theory while claiming to apply it
The authors review the Wigner’s Friend scenario and then comment:
“Now, if we think that measurement M must have a definite outcome, then clearly we must conclude that either Wigner or Bob is wrong. Wigner is wrong according to any interpretation which tells us that people cannot be in superpositions - e.g. gravitational collapse interpretations and spontaneous collapse interpretations - because those interpretations affirm that there will be an actual or effective collapse of the wavefunction at around the time of Bob’s measurement so that Bob will no longer be in a coherent superposition by the time of Wigner’s measurement.”
This is a very peculiar statement. The assertion “Wigner is wrong” arises only from the authors’ refusal to allow Wigner to apply the very OR formulation they claim to be applying. As discussed above, OR instructs Wigner not to describe Bob by a macroscopic superposition but by a proper mixed state. The authors even admit this: “those [OR] interpretations affirm that there will be an actual or effective collapse of the wavefunction at around the time of Bob’s measurement so that Bob will no longer be in a coherent superposition by the time of Wigner’s measurement.” So why do they assume that Wigner will describe Bob by a coherent superposition dictated not by OR but by CQM? Perhaps more to the point, why do the authors ignore the fact, just stated by them, that the OR theory dictates that “Bob will no longer be in a coherent superposition by the time of Wigner’s measurement”? The fact that OR theories preclude the WF-type paradoxes is acknowledged by Baumann and Brukner (2019, 3), who say:
“[OR] theories postulate a modification of standard quantum mechanics, such as spontaneous [7] and gravity-induced [4, 8] collapse models, which become significant at the macroscopic scale. These models could exclude superpositions of observers as required in Wigner’s friend-type experiments and hence the protocol cannot be run.”  [^6] Baumann and Brukner go on to say “In our view this is the most radical position as it would give rise to new physics..”. However, the widespread assumption that OR theories necessarily depart from the predictions of quantum theory is incorrect, since at least one OR-type theory, the Transactional Interpretation, not only does not depart from the Born Rule but physically derives it. So any “new physics” of TI is confined to the very old but overlooked physics of the direct-action theory of fields, which yields an empirically equivalent form of quantum theory that also accounts for “measurement” (cf. Kastner 2022, Chapter 5).
Now, if one thought a particular OR formulation failed to accomplish actual reduction at a macroscopic level, one could quibble with that. But that is not what AR2022 have done. They admit that an OR theory precludes macroscopic superpositions. Yet they unaccountably persist in assuming that the WF scenario still applies, falsely implying that the OR theory does not give a satisfying resolution of a ‘paradox’ to which it is not even subject, while their proposed interpretation supposedly does.
2b. Error 2: Failure to acknowledge empirically consequential nature of inconsistent predictions
The inconsistencies that arise in the MP-generated paradoxes involve supposedly observed facts or predictions that are in conflict among different observers. Many discussions of these inconsistencies treat them as “incommensurable realities.” I will term such claims IRC, the “incommensurate realities” claim, exemplified by AR2022 as above:
(IRC) “Wigner and Bob simply live within incommensurate realities now, and no attempt to reach across and bridge the gap can possibly succeed.”
AR then attempt to remedy that asserted situation with what they term ‘cross-perspective links’ (CPL) as proposed in Postulate 4.1. However, this revision does nothing to remedy the fact that the ‘inner observer’ can indeed find that his prediction deviates from that of the ‘outer observer’ (according to CQM). This is demonstrated in an extant counterexample to IRC (discussed in Kastner 2020, 2021, 2022, and Baumann and Brukner 2019). The following material is excerpted from Kastner 2022. The term ‘UOA’ stands for ‘unitary outcome assumption,’ which is the standard working assumption under CQM that outcomes (somehow) result from a unitary interaction alone (even though, as noted above, CQM fails to account for outcomes).
{addmargin}
[2em]2em
The common assumption that inconsistencies involving specific outcomes must always remain hidden, based on the idea that they are only observer-relative, is untenable via a specific counterexample. The example (similar to a scenario of Baumann and Brukner, 2019) demonstrates that attributing outcomes to entangled subsystems, even if designated as only “relative” to the subsystem or involving only “subjective collapse,” (as discussed by Baumann and Wolf, 2018), leads to empirical inconsistency. [^7] Deutsch (1985) considers a scenario with some similarities, but presents it as an empirical comparison between the Everettian picture and the Copenhagen interpretation. The present work considers implications of this type of empirical disagreement that do not seem to have yet been fully appreciated.
Assume, as in Proietti et al (2019), that microscopic quantum systems such as atoms, molecules, and photons can act as “observers” whose correlations with other similar degrees of freedom constitute “measurements” (this is the UOA). Let $W$ and $F$ comprise degrees of freedom subject to 2D state spaces, such as spin-$\frac{1}{2}$ atoms. Let us suppose that $F$ is “measuring” another spin-$\frac{1}{2}$ system $A$, prepared in equal superposition of outcomes ‘up’ and ‘down’ along a direction $Z$, via a Stern-Gerlach device. $F$ comprises 2 degrees of freedom: $B$, acting as a pointer/memory, and $C$, for communication. $C$’s possible states are ‘ground’ and ‘excited,’ where the latter would be triggered by a photon signal from $W$. While this is a thought experiment and does not depend on any particular experimental setup, one could think of $F$ as a micro-detector whose $B$ states are correlated with $A$ via an appropriate interaction Hamiltonian. $C$ remains in its initial unexcited state $|0\rangle$ at this stage. Thus, based on the idea that an outcome is relative to an observer, according to $F$ after the “measurement” of $A$ by $B$, $A$ the relevant degrees of freedom are either in the state:
(1)1( 1 ) $|\Psi\uparrow\rangle_{FA}=|\uparrow\rangle_{A}\odot|\uparrow\rangle_{B}\odot|0\rangle_{C}$
or
(2)2( 2 ) $|\Psi\downarrow\rangle_{FA}=|\downarrow\rangle_{A}\odot|\downarrow\rangle_{B}\odot|0\rangle_{C}$
with equal probability–i.e., they are in a mixed state. On the other hand, according to $W$, $A$ and $F$ end up in a pure Bell state with $F$’s communication degree of freedom $C$ along for the ride:
(3)3( 3 ) $|\Psi\rangle_{FA}=|\Phi^{+}\rangle|0\rangle_{C}=\frac{1}{\sqrt{2}}\bigl{(}|\uparrow\rangle_{A}|\uparrow\rangle_{B}+|\downarrow\rangle_{A}|\downarrow\rangle_{B}\bigl{)}\odot|0\rangle_{C}$
Now, let $W$ subject the $B+A$ system to a measurement of the Bell observable for which the state $|\Phi^{+}\rangle$ is an eigenstate. The Bell observable measurement is somewhat involved, but there is nothing to preclude it in principle. $W$ could comprise as many bilevel systems as one wishes in order to allow for a pointer, signaling system, and any other degrees of freedom needed to implement the required correlations via interaction Hamiltonians, under the assumption (UOA) that all that is required for a “measurement outcome” (or outcomes in an Everettian approach) is a one-to-one correlation between eigenstates of an observable of the system being measured and the states of the putative “observer” (as assumed in Proietti et al). [^8] The Bell observable is not uniquely required. Alternatively, one could arrange for $F$’s pointer $B$ to be anti-correlated with $A$’s putative outcomes, and have $W$ conduct a measurement of the total spin of the system. This would reveal the inconsistency since $W$ predicts that the total spin $J$ must always be found to be 1, while according to F, W has a 50% chance of finding the result $J=0$, for the singlet state. $W$’s measurement is accompanied by a signal to $F$ as follows. An outcome finding the state $|\Phi^{+}\rangle$ results in a photon being emitted to $F$ to excite his communication degree of freedom $C$. This could be implemented by having $W$’s outcome registration involve excitation of a degree of freedom with a very short decay time such that it immediately de-excites in a collimated direction to $F$. According to $W$, $F$ should receive that photon for every run of the experiment. This makes the inconsistency manifest, since according to F, his probability of receiving the photon is only 1/2 given his description of the situation by the mixed state (1,2). (This concludes material excerpted from Kastner, 2022.)
We should note that the ”basic postulate of RQM” (Rovelli 1986), repeated by Laudisa and Rovelli 2021) is refuted by this counterexample:
“The physical assumption at the basis of RQM is the following postulate: The probability distribution for (future) values of variables relative to $S$ depend on (past) values of variables relative to $S$ but not on (past) values of variables relative to another system $S^{\prime}$.”
Thus, RQM asserts that $F$ should be able to predict correctly the future values of its own local system(s) by reference only to its local past outcomes, while its presumed unitary entanglement with the result of the world continues. But in the counterexample (presuming that $W$ is correct, as AR assert), $F$ cannot correctly predict the probability of a future event relative to itself—i.e., the probability of its degree of freedom $C$ becoming excited—based on its own past outcomes. This probability apparently depends instead on another system outside ($W$), contrary to the central postulate of RQM. At this point, RQM is effectively dead. The version offered in AR2022, RQM22, is a mere shadow of its former self. Yet even that shadow fails to survive, as we shall see in what follows.
2c. Unsupported claim about ‘destroyed information’
The authors first focus on alleged incommensurability of the inner and outer observers with respect to a commuting observable, i.e. (with Alice playing the role of F and Bob playing the role of W). They say:
“Now suppose Bob measures S in the same basis as Alice’s measurement, and hence he obtains a measurement outcome MBS which he will interpret as providing information about the result of Alice’s measurement on S. Suppose that Bob also ‘measures’ Alice herself and obtains a measurement outcome MBA for the value of of some pointer variable which is supposed to be a record of her measurement result - for example, he could simply ask her what her measurement result was. So in this scenario we have three different measurement outcomes MA,MBS,MBA all supposedly providing information about the value of the same variable.” (AR2022)
The authors then lament the alleged fact that these results are incommensurable and go on to replace Postulate 4 with Postulate 4.1 concerning “Cross-perspective Links”, which we repeat here for convenience:
Definition 4.1. Cross-perspective links: In a scenario where some observer Alice measures a variable V of a system S, then provided that Alice does not undergo any interactions which destroy the information about V stored in Alice’s physical variables, if Bob subsequently measures the physical variable representing Alice’s information about the variable V, then Bob’s measurement result will match Alice’s measurement result.” (AR20220
The provision about destroyed information does a lot of ‘heavy lifting’ in the authors’ subsequent attempt to remove the threat to RQM of the Wigner’s Friend paradox (as we will discuss further below). However, the system in the counterexample has a degree of freedom C that is not subject to any interaction that would ‘destroy’ its information, and F makes a specific prediction about the probability of a state change in C that disagrees with the probability predicted by W. If one takes the ‘correct’ description as the that of F being in a macroscopic superposition, then F must find out he is wrong, violating the central postulate of RQM as noted above. Does RQM22 do any better? AR2022 offer the following:
“But when postulate four is replaced with cross-perspective links, we have a different story. In this approach, Bob performs his measurement and the system he measures takes on a definite value relative to him, and then perhaps he will be tempted to reason according to the steps set out in ref [42] to conclude that the measurement M will definitely have outcome 0. But relative to Wigner, Bob is still in a superposition state and thus Wigner is able to perform a measurement on Bob in which the value of the variable measured by M takes on the value 1 relative to Wigner. Moreover, although this value is relativized to Wigner, it is also an observer-independent fact [emphasis added] in the sense that other observers can find out about it by measuring or indeed just asking Wigner - for example, if Bob gets out of the box and asks Wigner what has transpired, Wigner will tell him that the measurement had outcome 1 and Bob will perceive Wigner saying that the measurement had outcome 1. So Bob is wrong, although since the measurement on Bob will destroy the information stored in Bob’s physical variables about the outcome of his measurement and hence also any inferences he made on the basis of that measurement, Bob will necessarily have no memory of making a wrong prediction, so no direct contradiction will ever arise.”
However, in the above counterexample, according to CQM (which assumes there is no real reduction), the inner observer will indeed find that he is ‘wrong’ just by checking his degree of freedom C. There is no ‘box’ to step out of, nor any mechanism for wiping his memory. The assumption that the inner observer’s memory will be wiped is merely an article of faith that is falsified by the counterexample. RQM22 fares no better than RQM, falling victim to CQM’s acknowledged inability to account for determinate outcomes.
2d. Infinite regress involving ‘superobservers’ nullifies consistency claim
As noted above, the authors assert that Wigner’s outcome is an ‘observer-independent fact’ (BF sentence). Yet CQM fails to license this assumption, since one could always have another ‘outside observer’ Wigner2 who (according to CQM) must describe Wigner as being in a coherent superposition after his measurement—since according to CQM, no observer is ever correct in thinking that any collapse to an eigenvalue actually occurred. This means that Wigner’s outcome cannot be an ‘observer independent fact’ (at least as a piece of information supporting a correct prediction) as asserted in the quoted excerpt. This point reveals the fallacy in the following argumentation by the authors:
“Note also that Bob’s erroneous prediction arises only because he assumes that there is some kind of collapse when he performs his measurement. If instead he uses the version of RQM that we have suggested in this article, then he will correctly predict that Wigner could still get the result 1, since he will be aware that although his observation has caused him to update his own relative state, he remains in a superposition state relative to Wigner (since that state describes the relational [sic] between Wigner and Bob, not the absolute state of Bob). So in fact no contradictory predictions will arise as long as Bob and Wigner are consistent about employing the same interpretation of quantum mechanics.” (AR2022)
The infinite regress becomes obvious here if we simply replace Bob by Wigner and Wigner by Wigner2:
“Note also that Wigner’s erroneous prediction arises only because he assumes that there is some kind of collapse when he performs his measurement. If instead he uses the version of RQM that we have suggested in this article, then he will correctly predict that Wigner2 could still get the result X, since he will be aware that although his observation has caused him to update his own relative state, he remains in a superposition state relative to Wigner2 (since that state describes the relational [sic] between Wigner2 and Wigner, not the absolute state of Wigner). So in fact no contradictory predictions will arise as long as Wigner and Wigner2 are consistent about employing the same interpretation of quantum mechanics.”
In other words, the ostensible remedy for the MP-related inconsistencies, according to RQM, is to assume that every observer’s prediction, based on his measurement outcome and assignment of the corresponding eigenstate, is potentially erroneous based on the possible existence of an ‘outer observer.’ Along with this goes the untenable assumption that the inner observer will have no memory of being wrong, which is falsified by the counterexample discussed above. Thus, RQM prescribes that all experimenters are wrong when they attribute outcome eigenstates to their systems, since some outside observer would deny that attribution. But there could never be any correct ‘outside observer’ due to the infinite regress of ‘outside observers’. And indeed there can be no ‘observer independent facts’ if all observers are mistaken whenever they think that they got an outcome corresponding to an eigenstate of the measured observable. RQM thus fails to support any observer-independent facts, contrary to the authors’ claim, i.e.:
“…we now regard the pointlike quantum events or ‘flashes’ as absolute, observer-independent facts about reality, rather than relativizing them to an observer.” (AR2022)
This simply will not do, since there is no physics given to support such a ‘quantum event’. If Bob makes a measurement and is taken as finding an outcome, the authors seem to want to take that as a ‘quantum event’ corresponding to a ‘fact’. But under CQM (and RQM and RQM22), that ‘fact’ must be wrong, and so is Wigner’s, ad infinitum. In what coherent sense could myriads of wrong impressions possibly count as ‘observer-independent facts about reality’?
A further note is in order regarding the approach advocated in Baumann and Brukner (2019) (henceforth BB2019), which seeks to tame the WF-type inconsistencies by recourse to an instrumentalist treatment. They consider essentially the same counterexample discussed in Section 2b, in which an ‘inner observer’ could (according to CQM) find out they were wrong about their outcome-based prediction. BB2019 proceed under the conventional assumption that CQM is the correct theory and that the resulting WF-type scenarios are arguments against the existence of ‘observer-independent facts’, as in their remark:
“In Ref. [2, 3] a combination of two spatially distant Wigner’s-friend setups is used as an argument against the idea of ‘observer-independent facts’ via the violation of a Bell-inequality.”
They then offer an instrumentalist approach involving ‘rational agents’ to attempt to accommodate the resulting inconsistencies, in which agents can update their ‘degree of belief’ subject to new information from an ‘outside observer’. However, BB2019 allow for an ‘automatic machine’ to serve as ‘Wigner’ and to obtain an outcome which it then passes to the Friend (to get around worries about Wigner being dishonest about his outcome). The authors appear not to notice that they have just helped themselves to an observer-independent fact, namely the machine’s outcome. In other words, in attributing an outcome to a machine, they tacitly assume that quantum theory does in fact apply to physical systems, not ‘agents’, so that a ‘fact’ can obtain without an agent. Which is it? Are the ‘facts’ corresponding to outcomes observer-dependent or not? The discussion in BB2019 assumes both, despite the fact that they are mutually exclusive. Moreover, in any case, CQM fails to predict any outcomes. If one helps oneself to an outcome despite that, the outcome (or at least any ‘fact’ it might be taken to imply) is subject to being wrong, whether it is attributed to an agent or to a physical system. And the latter has no ‘degree of belief’ to update. The theory—CQM—that is unaccountably taken as yielding an outcome (or ‘fact’) simply cannot provide a correct one.
3. Conclusion
It has been shown that a recent attempt to shield Relational Quantum Mechanics (RQM) from inconsistencies arising from conventional quantum theory’s assumed unitary-only character, and its attendant measurement problem, fails. The attempt founders on an explicit counterexample to the usual claim that discrepancies in predictions of ‘inner’ and ‘outer’ observers in Wigner’s Friend-type scenarios remain ‘hidden’ or incommensurable, as well as other errors: (1) failure to actually apply an objective reduction model when the authors claim to be applying it, (2) an unsupportable claim that the outer observer’s measurement will ‘destroy’ information available to the inner observer (refuted by the discussed counterexample), and (3) an overlooked infinite regress in which no observer can ever be considered correct because he or she might be an ‘inner observer’; there is no natural or consistent end-point for the designation of ‘outer observer.’ This situation also negates the authors’ claim that their account provides for ‘observer-independent facts.’
The above points should also be kept in mind when evaluating the preferred options of Baumann and Brukner (2019), who treat quantum theory in an instrumentalist fashion, as a tool for making predictions. They refer to the inner observer as being subject to ‘updating degrees of belief’, but as noted above, under that approach there is still in principle an infinite regress of observers whose ‘observed outcomes’ are neither predicted by CQM nor can be regarded as facts about reality. Yet even the latter treatment implicitly smuggles in the observer-independent occurrence of outcomes in assuming that an ‘automatic machine’ could play the part of the outer observer. In that case, the infinite regress applies, and the ‘automatic machine’ could somehow be ‘mistaken’ about its outcome. Thus, once again, CQM is faced with a fatal inconsistency among outcomes that no resort to ‘updating’ can cure, since a machine is not subject to ‘degrees of belief’.
In contrast to this unsatisfactory situation facing conventional quantum theory and its putative interpretation RQM, it has been recalled that theories that incorporate objective reduction are not subject to the measurement problem (as acknowledged by one of the authors in a separate publication, Adlam 2023) and thus are not subject to its attendant paradoxes that lead to inconsistencies subject to remedy only by resort to observer-dependent instrumentalism and ad hoc postulations of outcomes not predicted by the theory itself, and in which all such postulated outcomes are subject to yielding wrong predictions.